OpenStudy (anonymous):

Once canned Juice drink is 25% orange juice; another is 5%. How many liters of each should be mixed together in order to get 20L that is 22% orange Juice?

OpenStudy (anonymous):

25% orange juice; let it be x liters another is 5%orange juice. this will be 20-x liters 20L that is 19% orange juice ... 0.25x+0.05(20-x) = 0.19*20 0.25x+1-0.05x=1.8 0.20x=1.8-1 0.20x=0.8 x=4 liters of 25% juice 16 liters of 5% juice

OpenStudy (anonymous):

Ratio=\(17:3 \implies \) 17 L and 3 Liters of 25% and 5% respectively.

OpenStudy (anonymous):

Classic system of equations problems. Let x be the amount of the 25% juice drink; y will be the amount of 5% juice drink. We get two equations x + y = 20 .25x + .05y = .22(20)

OpenStudy (anonymous):

If you convert the second to fractions, and multiply by a convenient factor, the system can solve without too much pain, even without a calculator.

OpenStudy (anonymous):

This is your final warning @Rohangrr, Please stop copy pasting wrong answers from other source. Ref: http://www.algebra.com/algebra/homework/equations/Equations.faq.question.351486.html

OpenStudy (anonymous):

oh!!! @FoolForMath okay!!! hows that wrong!

OpenStudy (anonymous):

@lcombs: @Rohangrr 's answer is wrong for this question.

OpenStudy (anonymous):

@Rohangrr: "20L that is 19% orange juice"